A parallel plate capacitor is constructed of metal plates, each with an

area of 0.20 m2. The capacitance is 7.9x10-9 F. Determine the plate

separation distance.


1
Expert's answer
2022-05-24T13:56:25-0400
"C=\frac{\epsilon_0 A}{d}"

"d=\frac{\epsilon_0 A}{C}\\
d=\frac{8.85*10^{-12}*0.20}{7.9*10^{-9}}=2.2*10^{-4}\:\rm m"